# Coercivity Panorama of Dynamic Hysteresis

### v1

@@ Abstract
We end this Letter with three remarks: i) Currently, the scaling exponents associated with hysteresis loops measured experimentally in different magnetic systems vary, and there are ambiguities in their theoretical explanations (See End Matter for details). The coercivity panorama we present provides a reliable basis for experimentally measuring the dynamic scaling of hysteresis loops. In particular, the finite-size scaling of coercivity is expected to be verified using microscopic experimental platforms such as cold-atom systems [66, 51] or two-dimensional ferromagnetic materials [82, 83]. ii) The phenomenological models used in the field of magnetism to study hysteresis loops and the research on the dynamics of FOPT in statistical physics previously seemed to be two independent areas. This work establishes a connection between them and presents a theoretical framework for coercivity based on microscopic dynamics. iii) The dynamics of FOPT and DPT can be uniformly characterized by coercivity. On this basis, the finite-time thermodynamics of interacting systems with phase transition driven at different time scales can be systematically studied. The optimal control of finite-time phase transitions [80, 54] with minimal energy cost, the performance of thermal machines with phase transition [84, 85, 86], and the information thermodynamics of interacting systems [87] are ̵fascinating issues worthy of further exploration.

### v2

@@ Abstract
We end this Letter with three remarks: i) Currently, the scaling exponents associated with hysteresis loops measured experimentally in different dynamic systems vary, and there are ambiguities in their theoretical explanations (See End Matter for details). The coercivity panorama we present provides a reliable basis for experimentally measuring the dynamic scaling of hysteresis loops. In particular, the finite-size scaling of coercivity is expected to be verified using microscopic experimental platforms such as cold-atom systems [57, 47] or two-dimensional ferromagnetic materials [75, 76].
ii) The phenomenological models of hysteresis used in the realm of magnetism and the research on the dynamics of FOPT in statistical physics previously seemed to be two independent areas. This work establishes a connection between them and presents a theoretical framework for coercivity based on microscopic dynamics. Besides the coercivity itself, the susceptibility at the coercive point, $\chi_{c}\equiv\evaluated{d\langle\phi\rangle/dH}_{H=H_{c}}$, is another crucial characteristic quantity [77, 52, 78]. According to Eq. (6), the finite-time and finite-size scalings of $\chi_{c}$ can be naturally obtained from the coercivity landscape.
iii) The dynamics of FOPT and DPT can be uniformly characterized by coercivity. On this basis, the finite-time thermodynamics of interacting systems with phase transition driven at different timescales can be systematically studied. Previous studies [79, 80, 67, 68, 69] in the near-equilibrium regime have revealed that different driving protocols, although not affecting the $v_{H}$-scaling, can influence the irreversibility by changing the scaling coefficient. In this work, for simplicity, we use a linear field sweep to determine the finite-time scalings of the coercivity. Exploring the optimal control strategies for the coercivity and corresponding energy cost in regimes beyond $v_{H}$-scaling is of great significance for the optimization of phase transition systems [81, 50]. Besides, the performance of thermal machines with phase transition [82, 83, 84], and the information thermodynamics of interacting systems [85, 86] are ̵fascinating issues worthy of further exploration. Furthermore, given the observed dependence of the DPT’s driving regime on the set field amplitude, a detailed investigation based on coercivity behavior is particularly warranted [27, 33, 87].